OpenStudy (anonymous):

I need help check my answer. Please! Jesse has eight friends who have volunteered to help him with a school fundraiser. Five are boys and 3 are girls. If he randomly selects 3 friends to help him, find each probability.

OpenStudy (anonymous):

1. P(2 boys or 2 girls) =(2/8)+(2/8)=1/2 2. P(at least 2 girls) 2/8= 1/4 ?

OpenStudy (dumbcow):

ok to find these probabilities, divide number of favorable combinations over total possible combinations. There are 4 possible outcomes: GGG, GGB, GBB, BBB Since there are only 3 girls to choose from, the number of combinations are 3Cx where x is number of girls out of the 3. Since there are 5 boys to choose from, the number of combinations are 5Cx where x is number of boys out of the 3. \[P(GGG) = \frac{\left(\begin{matrix}3 \\ 3\end{matrix}\right)\left(\begin{matrix}5 \\ 0\end{matrix}\right)}{\left(\begin{matrix}8 \\ 3\end{matrix}\right)}\] \[P(GGB) = \frac{\left(\begin{matrix}3 \\ 2\end{matrix}\right)\left(\begin{matrix}5 \\ 1\end{matrix}\right)}{\left(\begin{matrix}8 \\ 3\end{matrix}\right)}\] \[P(GBB) = \frac{\left(\begin{matrix}3 \\ 1\end{matrix}\right)\left(\begin{matrix}5 \\ 2\end{matrix}\right)}{\left(\begin{matrix}8 \\ 3\end{matrix}\right)}\] \[P(BBB) = \frac{\left(\begin{matrix}3 \\ 0\end{matrix}\right)\left(\begin{matrix}5 \\ 3\end{matrix}\right)}{\left(\begin{matrix}8 \\ 3\end{matrix}\right)}\]

OpenStudy (dumbcow):

1. P(2 boys or 2 girls) =P(GBB) + P(GGB) 2. P(at least 2 girls) = P(GGB) + P(GGG)

OpenStudy (anonymous):

thank you for spend time answer my question

OpenStudy (dumbcow):

do you have any trouble doing the combinations?

OpenStudy (anonymous):

yes, I'm copnfusing , how I know Different between premutation and com bination, independent , dependent How I know probability add or mutiple

OpenStudy (anonymous):

I headache chapter 12

OpenStudy (anonymous):

I spend many time practice this chapter, I don't know how long take me to get it

OpenStudy (dumbcow):

don't worry, you'll get it. whenever they ask you to pick or select a group from a larger group, you need to think combinations or groups. This is the total number of groups and this is the number of groups they want me to find. Ask yourself these questions. Usually if they use "OR" it means add probabilities, if they use "AND" it means multiply